OpenStudy (scarlettfarra2000):

Please help someone Which arrangement for these fractions is from least to greatest 2/3, 3/5, 5/7 a) 3/5, 2/3, 5/7 b) 5/7, 2/3, 3/5 c) 2/3, 3/5, 5/7 d) 2/3, 5/7, 3/5

OpenStudy (anonymous):

2/3 = .66, 3/5 = .60, and 5/7 = .71. That should help.